If the vertices of a quadrilateral is given by (x2 – 4)2 + (y2 – 9)2 = 0 then area of quadrilateral is-
(A) 36 (B) 24
(C) 16 (D) 81

Grade:11

5 Answers

downey zer
6 years ago
here we see the sum of two squares is equal to zero
since both terms are positive and equal to 0
therefore both have to be equal to 0
therefore x=2,-2 and y=3,-3
area =24
hiphop aadhi2001
6 years ago
Downey Zer 

         Please Can you explain the steps clearly....???
What comes after this step?  How did you got the answer
therefore x=2,-2 and y=3,-3
area =24

 
siddhi
6 years ago
you will get after solving A =2,3 B=2,-3 C=-2,3 D =-2,-3 Now u can use the determinant method to find the answer remember to take the vertices in order .Thank u
downey zer
6 years ago
sorry for
downey zer
6 years ago
sorry for the delay. siddhi is right u can use the determinant method. OR what u can do is plot the points in the x-y plane and u will find that the points form a rectangle of length 4 units on one side and 6 units on other side. So i was able to tell the area to be 24 just by looking at the points. with enough practice u will also be able to find cool tricks to solve the problems easily. hope that helps.